On June 30, 2017, Wisconsin, Inc., issued $267,350 in debt and 18,400 new shares of its $10 par value stock to Badger Company owners in exchange for all of the outstanding shares of that company. Wisconsin shares had a fair value of $40 per share. Prior to the combination, the financial statements for Wisconsin and Badger for the six-month period ending June 30, 2017, were as follows: Wisconsin Badger $ (339,000) Revenues $(985, 000) 720,000 201,000 $ 138,000) (208, 000) (138,000) Expenses Net income Retained earnings, 1/1 Met income Dividends declared $ (265,000) $ (843,000) (265, 000) $ 106,250 Retained earnings, 6/30 $ (1,001,750) $ 346, 000) $ 110,750 433,000 929,000 727,000 $ 2,199,750 $59,000 Receivables and inventorv Patented technology (net) Equipment (net) 180,000 372,000 619,000 1,230,000 Total assets Liabilities Common stock Additional paid-in capital Retained earnings $ (568,000) (360, 000) (270,000) $ (414, 000) (200,000) (270,000) (1,001,750 346,000) $(1,230, 000) Total 1iabilities and equities $(2,199,750) Wisconsin also paid $30,800 to a broker for arranging the transaction. In addition, Wisconsin paid $43,100 in stock issuance costs. Badgers equipment was actually worth $765,250, but its patented technology was valued at only $350,700. What are the consolidated balances for the following accounts? (Input all amounts as positive values) Accounts a. Net income. Retained earnings, b. d. Goodwill f Common stock. g. Additional paid-in capital
Answer #1
Net income =265000(Wisconsin) - 30800 234,200
Retained earnings 1.1.17 843000
Patented technology =929000+350700 1,279,700
Goodwill (see working note below)
Liabilities =568000+414000+267350(debt) 1,249,350
Common stock =360000+(18400*10) 544,000
Add: paid-in capital =27000+(18400*(40-10))-43100(issuance cost) 535,900
Working note:Calculation of Goodwill

Consideration transfered (Face value ) for Badger stock

=(18400*40) +267350(debt)

1,003,350
Less:Book value of Badger=1230000(total assets -414000(liabilities) 816,000
Fair value in excess of book value 187,350
Excess fair value (under valued equipment) =765250-619000 146,250
Excess fair value (over valued Patented technology) =372000-350700 (21,300)
Goodwill =187350-146250-(-)21300 62,400

IN CASE OF ANY DOUBTS OR CORRECTIONS FEEL FREE TO COMMENT BELOW. PLEASE RATE MY ANSWER BY HITTING ??. THANK YOU